# Asymptotic geometry of non-abelian Hodge theory and Riemann–Hilbert correspondence, rank three $\widetilde{E}_{6}$ case

### v1

@@ 6.2.6 Phases of the trace coordinates in terms of holonomy of the spectral line bundle
In the sequel we will use the notation of the holonomy isomorphism (12).
@@
    $\frac{X/Z}{|X/Z|}\approx\operatorname{hol}_{B}(\mathcal{L}),\quad\frac{Y/Z}{|Y/Z|}\approx\operatorname{hol}_{2B}(\mathcal{L})$
@@
    $\frac{X/Z}{|X/Z|}\approx\operatorname{hol}_{A+3B}(\mathcal{L}),\quad\frac{Y/Z}{|Y/Z|}\approx\operatorname{hol}_{2B}(\mathcal{L})$
@@
    $\frac{X/Z}{|X/Z|}\approx\operatorname{hol}_{A}(\mathcal{L}),\quad\frac{Y/Z}{|Y/Z|}\approx\operatorname{hol}_{-B}(\mathcal{L})$
@@
    $\frac{X/Z}{|X/Z|}\approx\operatorname{hol}_{A}(\mathcal{L}),\quad\frac{Y/Z}{|Y/Z|}\approx\operatorname{hol}_{2A}(\mathcal{L})$
@@
We only prove the first statement of the first item and leave the rest to the readers, because the proofs follow the same line. We apply formulas (28) combined with Proposition 6.9. On $S_{2}$, the dominant term of $X,Y,Z$ are respectively their first, second and first terms. The effect of the recessive terms on the angles is negligible, and they only amount to a homotopy. We find
$\sqrt{-1}\operatorname{Im}(\ln(X/Z))\approx\int_{\eta_{0}-\eta_{1}}u_{2}-u_{3}+\int_{\gamma_{0}-\gamma_{1}}u_{2}$
@@ 6.2.7 Asymptotics over the Stokes rays
    including the possibility that the latter two limits are both $\infty$ (but only simultaneously).
@@
In particular, this dependence is continuous. In order to find desired limits, we need the following statement.
Lemma 6.15.
The asymptotic expansions of (28) hold over the Stokes rays too.
In view of Proposition 6.4, this would follow from the following more technical statement, applied individually to all diagonal entries appearing in $X_{R,\varphi},Y_{R,\varphi},Z_{R,\varphi}$ (up to replacing $\infty$ by 00 and setting $r=R^{-1}$).
Lemma 6.16.
Let
[…]
be an $\mathbb{R}$-analytic function. Assume that for each $\varphi\in(0,2\pi)\setminus\{\pi\}$ we have that
$f\left(re^{\sqrt{-1}\varphi}\right)\to 1\quad as\;r\to 0.$
Then we also have
$f(\pm r)\to 1\quad as\;r\to 0.$
This property should be a consequence of the basic rigidity of analytic functions, however as yet we lack an appropriate reference or a counterexample.
Using the above result gives that the maximal dilation exponents of the trace coordinates agree:
@@ 6.3.1 Critical angles and arc decomposition
6.3.2 Simpson’s map
Now consider the blow up of the compactifying curve at its nodal point. Locally around the nodal point, the curve $C$ consists of two components, namely the $X$ and $Y$ axes in the local coordinate system $[X\colon Y\colon Z]$. After the blow up at its nodal point $[0\colon 0\colon 1]$, the proper transform of $C$ is denoted by $C_{1}$, and the exceptional curve by $C_{2}$, so that $C_{1}$ and $C_{2}$ have two intersection points. Denote the intersection point of $C_{2}$ with the proper transform of the $X$ axis by $P_{1}$, and the one with the $Y$ axis by $P_{2}$. The component $C_{i}$ gives rise to the vertex $V_{i}$ in the dual graph, and the two intersection points $P_{1},P_{2}$ correspond to edges $E_{1}$ and $E_{2}$ respectively. Then the dual boundary graph becomes as on Figure 6.
$S\colon\mathcal{M}_{B}\setminus\psi(h^{-1}(B_{R}(0)))\to\lvert\mathcal{D}\partial\mathcal{M}_{B}\rvert$ (33)
as follows. We consider the compactification ¯$\overline{\mathcal{M}}_{B}$ of $\mathcal{M}_{B}$ as the blow-up of its projectivization at the nodal point. Next, we construct a cover
$\mathcal{M}_{B}\setminus\psi(h^{-1}(B_{R}(0)))=N_{0}^{*}\cup N_{1}^{*}\cup N_{2}^{*}$ (34)
  1. 1.
    $N_{i}^{*}\cap N_{j}^{*}$ is non-empty and connected for every $0\leq i,j\leq 2$
  2. 2.
    $N_{0}^{*}\cap N_{1}^{*}\cap N_{2}^{*}=\emptyset$.
Namely, let $N_{0}$ be a tubular neighborhood of $C_{1}$ in ¯$\overline{\mathcal{M}}_{B}$ and $N_{1},N_{2}$ be suitable neighborhoods of the intersection points $P_{1},P_{2}$ of $C_{1},C_{2}$ defined above. Let
$N_{i}^{*}=N_{i}\cap\mathcal{M}_{B}=N_{i}\setminus(C_{1}\cup C_{2}).$
$\phi_{1}+\phi_{2}+\phi_{3}=1.$
Its edges are defined by the $\phi_{i}=0$ plane sections.
@@ 6.3.3 Identifying the images of suitable arcs
Connectedness of the image follows in each case by continuity. A natural pair of homogeneous coordinates $[X^{\prime}\colon Y^{\prime}]$ on the exceptional curve $C_{2}$ satisfies
$\frac{X}{Y}=\frac{Y^{\prime}}{X^{\prime}}.$
@@
$[X\colon Y\colon Z]$ $\in N_{0}\cap N_{2}\cap C_{2}\Rightarrow\left|\frac{Y^{\prime}}{X^{\prime}}\right|<\frac{1}{10}$ (35)
$[X\colon Y\colon Z]$ $\in N_{0}\cap N_{1}\cap C_{2}\Rightarrow\left|\frac{X^{\prime}}{Y^{\prime}}\right|<\frac{1}{10}.$ (36)
@@
    According to Lemma 6.17, $\psi\circ\sigma$ maps $A_{2}\subset\operatorname{Int}I_{2}$ into a tiny neighborhood of the point $P_{2}$. Since $P_{2}\in N_{0}\cap N_{2}$ and the triple intersection of the $N_{i}$ vanish, we see that $P_{1}\notin N_{1}$. The statement follows because the partition of unity is subordinate to (34).

### v2

@@ 6.3.4 Phases of the trace coordinates in terms of holonomy of the spectral line bundle
In Lemma 6.12, we determined the dominance properties of the absolute values of the trace coordinates along rays in any open sector $S_{j}$. We have found that $X/Z\to 0$ and $Y/Z\to 0$ as $R\to\infty$. Moreover, in sectors $S_{j}$ with $j\equiv 2,3\pmod{4}$ the limit in $\widetilde{\mathcal{M}_{\operatorname{B}}}$ is $P_{2}$ and in the remaining ones it is $P_{1}$. Here, we focus on the angles of the affine coordinates near these limit points found in Section 6.3.1. We find that these angles can be interpreted as the holonomy of $\mathcal{L}$ along some loops in $\widetilde{\Sigma}$, depending on the given sector; as a matter of fact, the computation on the other sectors is the same up to a permutation of the sheets. Due to the $32\pi/$ 3 periodicity of the diagram of Figure 3, it is sufficient to deal with four consecutive sectors. In the sequel we will use the notation of the holonomy isomorphism (12).
@@
    $\frac{Y/X}{|Y/X|}\approx\operatorname{hol}_{B}(\mathcal{L}),\quad\frac{Y/Z}{|Y/Z|}\approx\operatorname{hol}_{2B}(\mathcal{L})$
@@
    $\frac{Y/X}{|Y/X|}\approx\operatorname{hol}_{-A-B}(\mathcal{L}),\quad\frac{Y/Z}{|Y/Z|}\approx\operatorname{hol}_{2B}(\mathcal{L})$
@@
    $\frac{X/Z}{|X/Z|}\approx\operatorname{hol}_{A}(\mathcal{L}),\quad\frac{X/Y}{|X/Y|}\approx\operatorname{hol}_{A+B}(\mathcal{L})$
@@
    $\frac{X/Z}{|X/Z|}\approx\operatorname{hol}_{A}(\mathcal{L}),\quad\frac{X/Y}{|X/Y|}\approx\operatorname{hol}_{-A}(\mathcal{L}).$
@@
We only prove the first statement of the first item and leave the rest to the readers, because the proofs follow the same line. We apply formulas (28) combined with Proposition 6.11. On $S_{2}$, the dominant term of $X,Y,Z$ are respectively their first, second and first terms. The effect of the sub-leading terms on the angles is negligible, and they only amount to a homotopy. We find
$\sqrt{-1}\operatorname{Im}(\ln(Y/X))\approx\int_{\eta_{0}-\eta_{1}}u_{2}-u_{3}+\int_{\gamma_{0}-\gamma_{1}}u_{2}$
@@
6.3.5 End of the proof
The homotopy class of a fiber of $S$ is given by the torus $T_{\varepsilon_{1},\varepsilon_{2}}$ of the form
$\left|\frac{X}{Z}\right|=\varepsilon_{1},\quad\left|\frac{X}{Y}\right|=\varepsilon_{2}$
in $U_{1}$ for some (equivalently, any) $0<\varepsilon_{1},\varepsilon_{2}\ll 1$. Now, if we choose $z_{1}\in S_{4}$ of absolute value $1$ then the image of $H^{-1}(Rz_{1})$ by $\psi$ lies in a neighbourhood of $T_{\varepsilon_{1},\varepsilon_{2}}$ for suitable values of $\varepsilon_{1},\varepsilon_{2}$. Moreover, by Proposition 6.15, the angles of the above affine coordinates are approximated by the holonomy of the spectral sheaf along $A,A+B$. The sub-leading terms of the asymptotic expressions only amount to a homotopy. As $A,A+B$ is a $\mathbb{Z}$-basis of $H_{1}(\widetilde{\Sigma},\mathbb{Z})$, we get the assertion.
6.4 Conditional proof of Proposition 6.3
@@ 6.4.1 Asymptotics over the Stokes rays
The results of this section are conditional on the following hypothesis, for which we currently have no proof.
Assumption 2.
The asymptotic expansions of (28) hold over the Stokes rays too.
However, we emphasize that the results of this section are not needed anywhere in the sequel, so the main results of the paper do not rely on this assumption. The only reason we include this section is to shed a light on the possible behaviour of the correspondence in the Stokes directions, which is otherwise not immediate to guess.
@@
    including the possibility that the latter two limits are both simultaneously $\infty$.
@@
Using Assumption 2 we see that the maximal dilation exponents of the trace coordinates agree:
@@ 6.4.3 Identifying the images of suitable arcs
Connectedness of the image follows in each case by continuity. Recall from Section 6.3.1 the natural pair of homogeneous coordinates $[X^{\prime}\colon Y^{\prime}]$ on the exceptional curve $C_{2}$ satisfying
$\frac{X}{Y}=\frac{X^{\prime}}{Y^{\prime}}.$
@@
$[X\colon Y\colon Z]$ $\in N_{0}\cap N_{2}\cap C_{2}\Rightarrow\left|\frac{X^{\prime}}{Y^{\prime}}\right|<\frac{1}{10}$ (35)
$[X\colon Y\colon Z]$ $\in N_{0}\cap N_{1}\cap C_{2}\Rightarrow\left|\frac{Y^{\prime}}{X^{\prime}}\right|<\frac{1}{10}.$ (36)
@@
    According to Lemma 6.17, $\psi\circ\sigma$ maps $A_{2}\subset\operatorname{Int}I_{2}$ into a tiny neighborhood of the point $P_{2}$. Since $P_{2}\in N_{0}\cap N_{2}$ and the triple intersection of the $N_{i}$ vanish, we see that $P_{2}\notin N_{1}$. The statement follows because the partition of unity is subordinate to (34).